OpenStudy (anonymous):

MEDAL AND FAN!!!! Which of the following represents the zeros of f(x) = 3x3 − 10x2 − 81x + 28

OpenStudy (usukidoll):

\[f(x) = 3x^3-10x^2-81x+28 \] this?

OpenStudy (usukidoll):

if we need to find 0's then we need the rational root test.

OpenStudy (usukidoll):

The Rational Roots Test (also known as Rational Zeros Theorem) allows us to find all possible rational roots of a polynomial. Suppose "a" is root of the polynomial P(x) that means P(a) = 0. In other words, if we substitute "a" into the polynomial P(x) and get zero it means that the input value is a root. so we take the last term, 28, and find all factors of 28 which is 1,2,4,7,14,28. We also have to take the negative versions as well. So, 1,-1,2,-2,4,-4,7,-7,14,-14,28,-28 just plug these numbers into the equation and stop once you have a 0 there could be more than one root, just one, or none. If we don't have any roots, the test fails.

OpenStudy (anonymous):

is the answer 7, -4, 1/3

OpenStudy (alekos):

but the last root lies somewhere between 0 and 1 and I think 1/3 is pretty close

OpenStudy (anonymous):

answer choices 7,-4,1/3 7,-4,-1/3 7,4,1/3 7,4,-1/3

OpenStudy (usukidoll):

hmmm.... let x = 7,4, or -4 and plug it into the equation ... one of them has to produce a zero .we may have to go further like long division to grab the remaining roots

OpenStudy (anonymous):

im pretty sure it is the last choice

OpenStudy (usukidoll):

did you test it out?

OpenStudy (anonymous):

kind of

OpenStudy (usukidoll):

because there's a negative root.. and that number is not -1/3

OpenStudy (usukidoll):

so that knocks out two choices.

OpenStudy (anonymous):

that means it is the third choice because after you test it out it gives you the third choice

OpenStudy (usukidoll):

no

OpenStudy (alekos):

negative root is x=-4

OpenStudy (usukidoll):

mhm ^

OpenStudy (anonymous):

oops i think i know the answer for this question can you help me with another

OpenStudy (usukidoll):

yeah sure. you actually had the right answer 10 minutes ago XD

OpenStudy (anonymous):

Which of the following is a polynomial with roots 5, 4i, and −4i

OpenStudy (anonymous):

f(x) = x^3-5x^2+20x-16 f(x) = x^3-5x^2+16x-80 f(x)=x^3-20x^2+5x-16 f(x) = x^3-16x^2+80x-5

OpenStudy (usukidoll):

wow... it's like we have to use the roots given and plugging in it into the function having x = 5 will make it easier though.

OpenStudy (anonymous):

so plug 5 for each one

OpenStudy (usukidoll):

try let x = 5 for the second function

OpenStudy (usukidoll):

yeah for each one, but I think I've weeded it out x=5 for the second function

OpenStudy (anonymous):

i think i can take out the 2nd one and also the the first one

OpenStudy (usukidoll):

first one is nasty... x.x

OpenStudy (anonymous):

so that is definitely eliminated

OpenStudy (usukidoll):

oh yeah... not to mention that the real root is x=1... not what we need.

OpenStudy (anonymous):

the answer is the third one

OpenStudy (usukidoll):

umm are you sure?

OpenStudy (anonymous):

im pretty sure

OpenStudy (usukidoll):

the real root I'm getting for that third choice is really bad ... same with the imaginary... the real root is approx.. 20.. we need x = 5,4i,-4i cross that out

OpenStudy (anonymous):

im feeling that my work shows that it is the answer

OpenStudy (anonymous):

the last one

OpenStudy (usukidoll):

try plugging x = 5 for the second function you have and tell me if your result is 0 f(x) = x^3-5x^2+16x-80

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

then that's the function. since the result became 0... for x = 5... then the roots x =4i, -4i work as well.

OpenStudy (anonymous):

thanks

OpenStudy (usukidoll):

you're welcome :)